Example of Phenomenological Viewing

Greg Hudson

October 7, 1994

Consider the example of viewing discarding a piece of trash in a waste basket (since, as philosophers, we must be willing to dredge up the dirt and garbage of the world). First, we note a relationship of instance between garbage cans in general and the actual garbage can into which we see someone discarding the piece of trash. Assuming we are familiar with the notion of garbage cans, we will normally see the garbage can (in the natural attitude) as the general garbage can, not as an instance. (If we are not familiar with garbage cans, we will see it as an instance of a general function which we do not know, and perhaps perform eidetic variation in order to discover what that function is.)

Second, we note layers of functionality of the garbage can. If we are viewing someone discarding a piece of trash and we are engaged in the project of wanting to littering, we may see the garbage can as a place for putting trash so it doesn’t get all over the floor. If we are engaged in the project of wanting to encourage recycling, we may see the garbage can as a place where people put paper when they should be recycling it.

We can also observe similar things about the piece of trash being discarded, but I won’t repeat myself. Instead, note how this example illustrates the conditions of possibility of viewing: we thematize the unthematically copresent feature that the garbage can is an instance of garbage cans in general, and therefore disclose a context (that the viewer understands garbage cans) which allows the situation to make sense. It is the function of the garbage can which gives it its role in the example (note that I did not even bother to stop and say that it was made of metal, or plastic, yet you understood what I was talking about). The natural attitude (“I see someone throwing a piece of trash in the garbage can”) is situated in the context of the scene–it is the attitude of a person who understands trash cans.and has a motivation for viewing the scene. As mentioned above, the project we are engaged in when we view the scene affects at which level we see the layered functionality of the garbage can and trash.